# references.bib

### v1

@@ 3.2 Models
[…]
Deep neural network models. Additionally, we consider different deep neural network architectures used for vision (convolutional networks, ResNets, and Vision Transformers, Section 4.4) and ReLU networks trained through backpropagation (Section 4.5).
Figure 1: a, Theoretical setup: we consider a disentangled representation of the input, followed by a nonlinear transform $\phi$ and a learned linear readout $w$. b, Training sets for symbolic addition. The grid represents nine components with associated values $-4,-3,\dotsc,4$. Each tile represents a data point in the training set. c, Context dependence. In context 1, feat. 1 determines the category; in context 2, feat. 2 determines the category. The training sets leave out different subsets of the lower right orthant. d, Transitive equivalence: six items are split up into two arbitrary equivalence classes (e.g. A,B,C and D,E,F) and generalization requires transitive inference over equivalence classes.
3.3 Example tasks
In addition to deriving a general theory, we consider a set of example tasks that are important building blocks for compositional reasoning in machine learning and cognitive science (Appendix D discusses additional tasks including logical operations and partial exposure):
Symbolic addition. Many tasks involve inferring a magnitude associated with different underlying components (e.g. handwritten digits) [lorenzi_numerosities_2021, sheahan_neural_2021]. We consider two components $(z_{1},z_{2})$ with unobserved assigned values $v_{1}(I_{1})$ and $v_{2}(I_{2})$. The target is the sum of those values: $y=v_{1}(z_{1})+v_{2}(z_{2})$. After sufficient exposition to individual items, a model with an additive structure can generalize to novel combinations of items [wiedemer_compositional_2023]. In particular, we consider nine input elements $[-4],[-3],\dotsc,[4]$ with associated values $-4,-3,\dotsc,4$ and training sets containing all pairs where at least one component is equal to a certain subset of values. This can require interpolation, extrapolation, or both (Fig. 1b).
Context dependence. The relevance of different stimuli often depends on the context we are in. Taking into account this context is a crucial aspect of cognition in humans and animals [bouton_stimulus_1999, taylor_context-dependent_2013, parker_cognitive_2014, ito_compositional_2022]. We therefore consider a task with three input components $(z_{\text{co}},z_{f1},z_{f2})$. The context ($z_{\text{co}}$) has two possible values specifying whether $z_{f1}$ or $z_{f2}$ determine the response. Both features have six possible values which are split up in two categories. If the model has learned this context dependence, it should be able to generalize to novel feature combinations. We evaluate on the orthant for which $z_{f1}$ indicates Cat. 2 and $z_{f2}$ indicates Cat. 1. In the most extreme generalization test (CD-3), we leave out the entire orthant, in easier versions we leave out conjunctions of two or one of those features (CD-2 and CD-1) (Fig. 1c).
Transitive equivalence. Relational reasoning is an important instance of compositional generalization and often involves extending the learned relations to new item combinations [halford_relational_2010, battaglia_relational_2018]. Given an unobserved (and arbitrary) equivalence relation, the task is to determine whether two presented items $(z_{1},z_{2})$ are equivalent. The model should generalize to novel item pairs using transitivity ($A=B$ and $B=C$ imply $A=C$) (Fig. 1d). This is an important instance of relational cognition (often studied as “associative inference” in cognitive science [schlichting_memory_2015, spalding_ventromedial_2018]). Note that prior work has found that kernel models often successfully generalize a transitive ordering relation [lippl_mathematical_2024].
4 Results
4.1 Kernel models with compositional structure are conjunction-wise additive
Our primary theoretical contribution in this work is to characterize the full range of compositional computations that can be implemented by kernel models with disentangled inputs: specifically, they assign a value to each combination of components seen during training and generalize to test inputs by adding up the relevant combinations’ values. We call this motif “conjunction-wise additivity.” Below we formally state our finding and then explain its implications.
We first note that the kernel of the disentangled representation, $K(z,z^{\prime})=z^{T}z^{\prime}$, only depends on the components for which the two inputs are overlapping, $O(z,z^{\prime}):=\left\{c|z_{c}=z_{c}^{\prime}\right\}$ where $c\in\{1,\dotsc,C\}$. This is necessitated by the fact that separate components are represented in orthogonal subspaces. We call any representation satisfying this criterion “compositionally structured.” Notably, the hidden representation $\phi(z)$ is also compositionally structured (in the infinite-width limit). This is because the kernel of the neural network’s representation, $K_{\phi}(z,z^{\prime}):=\phi(z)^{T}\phi(z)$, only depends on the inputs $z,z^{\prime}$ through their kernel $K(z,z^{\prime})$ and thus conserves this property [[]; see Appendix B]cho_kernel_2009,tsuchida_invariance_2018,tsuchida_richer_2019,han_fast_2022.
Intriguingly, we find that any kernel model with a compositionally structured representation is constrained to be conjunction-wise additive. To state this finding, we define, for each $z\in Z$, the set of conjunctions for which $z$ overlaps with some element in the training set $z^{\text{tr}}\in Z^{\text{train}}$,
$O(z|Z^{\text{train}}):=\left\{S\subseteq\{1,\dotsc,C\}|\exists z^{\text{tr}}\in Z^{\text{train}}\text{ s.t. }S\subseteq O(z,z^{\text{tr}})\right\}.$ (3)
We then prove:
Theorem 4.1.
For $z\in Z$, any kernel model $f$ with a compositionally structured representation and training data $Z^{\text{train}}$ can be expressed as a sum of functions $f_{S}$ over all conjunctions $S\in O(z|Z^{\text{train}})$, where each $f_{S}$ associates a value with each unique combination of features $z_{c}$ for $c\in S$:
$f(z)=\sum_{S\in\mathcal{O}(z|Z^{\text{train}})}f_{S}(z_{S}),\quad z_{S}:=(z_{c})_{c\in S}.$ (4)
We prove the theorem in Appendix A. Intuitively, it holds because the model associates a distinct weight vector with each conjunction/combination of components. The weight vectors associated with conjunctions not seen during training remain at their initial value and can therefore not be used.
To understand the theorem’s implications, we first consider a direct linear readout, $f(z)=w^{T}z$. This constrains the model to adding up a value for each component (“component-wise additivity”):
Component-wise additivity is a strict subset of conjunction-wise additivity (each $f_{S}$ depending on more than one component is set to zero). It lets the model generalize perfectly on component-wise additive tasks like symbolic addition. However, it also prevents the model from learning any training set that is not component-wise additive, including context dependence and transitive equivalence.
The nonlinear transform $\phi$ can overcome this constraint [hornik_multilayer_1989, cybenko_approximation_1989, leshno_multilayer_1993, rigotti_importance_2013]. Indeed, conjunction-wise additive functions more broadly can learn arbitrary training data, as the full conjunction $S=\{1,\dotsc,C\}$ is in $O(z|Z^{\text{train}})$ for each $z\in Z^{\text{train}}$ and $f_{S}(z)$ can take on a distinct value for each $z$. However, for $z\in Z^{\text{test}}$ the full conjunction is not in $O(z|Z^{\text{train}})$. Thus, while conjunction-wise additivity (unlike component-wise additivity) does not impose a constraint on the kinds of training data the model can learn, it does impose constraints on the kinds of generalizations it can implement.
In particular, for inputs with two components, $f(z)=f_{1}(z_{1})+f_{2}(z_{2})+f_{12}(z)$ for $z\in Z^{\text{train}}$. However, for $z\in Z^{\text{test}}$, $f_{12}(z)$ falls away and kernel models are constrained to an additive computation: $f(z)=f_{1}(z_{1})+f_{2}(z_{2})$. Thus, they cannot generalize on any non-additive task, including transitive equivalence: while they learn the training data (using $f_{12}(z)$), they cannot generalize to the test data.
For inputs with more than two components, a conjunction-wise additive computation does not just encode each component and their full conjunction, but also partial conjunctions between components (e.g. 12, 13, and 23 for three components). For test inputs, the model can still not rely on the full conjunction (123), but it can rely on any partial conjunction it may have seen before.
To see whether such a model can, in principle, generalize on a given task, we must 1) identify the overlaps with the training set, $\mathcal{O}(z|Z^{\text{train}})$, and 2) determine whether the target can be written as a conjunction-wise sum. For example, for context dependence, for all new test inputs, we have seen the context-feature conjunctions (i.e. 12 and 13), but not the feature-feature conjunction (i.e. 23). As a result, model behavior on the test set can be written as $f(z)=f_{1}(z_{\text{co}})+f_{2}(z_{f1})+f_{3}(z_{f2})+f_{12}(z_{\text{co}},z_{f1})+f_{13}(z_{\text{co}},z_{f2})$. While the task is not component-wise additive, we can use $f_{12}$ and $f_{13}$ to encode the task: $f_{12}(z_{\text{co}},z_{f1})$ encodes the target when the context $z_{\text{co}}$ indicates the first feature $z_{f1}$ as relevant and is zero otherwise; $f_{13}(z_{\text{co}},z_{f2})$ works in the opposite way.
Hence, conjunction-wise additivity not only tells us whether kernel models with compositionally structured representations can solve a certain task, but also how they solve it. Further, it tells us to make a task non-additive: in Section D.2.3, we describe a modification that makes context dependence non-additive by requiring generalizations to novel context-feature conjunctions.
Importantly, conjunction-wise additivity only determines whether certain compositional generalizations are, in principle, feasible. Model generalization also depends on whether the model identifies the correct conjunction-wise function. We discuss this in the next two sections.
Model generalization is determined by how prominently different component conjunctions are represented. This is measured by the similarity $K(S)$ between inputs with overlap $S\subseteq\{1,\dotsc,C\}$. $K(S)$ depends on responses to that particular conjunction $S$ as well as all of its sub-conjunctions $S^{\prime}\subsetneq S$. We thus define the scalar salience of an overlap $S$ as its unique contribution. We then normalize this salience, so all saliences add up to one:
Definition 4.2.
For a representation with kernel $K$ and a conjunction $S\subseteq\{1,\dotsc,C\}$, we define
$\overline{\text{Sal}}(\emptyset):=K(\emptyset),\quad\overline{\text{Sal}}(S):=K(S)-\sum_{S^{\prime}\subsetneq S}\overline{\text{Sal}}(S^{\prime}),\quad{\text{Sal}}(S)=\tfrac{\overline{\text{Sal}}(S)}{\sum_{\emptyset\neq S\subseteq\{1,\dotsc,C\}}\overline{\text{Sal}}(S)}.$ (6)
Figure 2: Normalized salience (i.e. $\text{Sal}_{C}(k)$) for neural networks with random weights and different nonlinearities for inputs with three components.
We now analyze how the representational salience evolves over different layers of a random network, considering an uncorrelated disentangled correlation with equal magnitudes (i.e. $\rho_{c}=0$ and $m_{c}=m_{c^{\prime}}$) (see Section B.2 for analysis beyond these assumptions). In that case, the salience only depends on the overlap size $k:=|S|$ and we denote it by $\text{Sal}_{C}(k)$.
Notably, conjunctions are only encoded and can only be used if they have non-zero salience. In particular, the disentangled representation has zero salience for all conjunctions of more than one component. As the network gets deeper, the salience of these conjunctions increases. In fact, for (leaky) ReLU networks, the salience of the full conjunction converges to one:
Proposition 4.3.
For a random neural network with a (leaky) ReLU nonlinearity, as $L\to\infty$, $\text{Sal}_{C}(k)\to 0$ for $k<C$ and $\text{Sal}_{C}(C)\to 1$.
We prove this statement in Section B.1. Empirically, we find that $\text{Sal}_{C}(C)$ indeed increases with depth, whereas $\text{Sal}_{C}(1)$ decreases (Fig. 2). The salience of intermediate conjunctions (e.g. $\text{Sal}_{3}(2)$) first increases and then decreases, its trajectory depending on the nonlinearity.
4.3 Kernel models suffer from two failure modes: memorization leak and shortcut bias
In Section 4.1, we found that kernel models sum up values assigned to different component conjunctions. This leaves ambiguous, however, whether they generalize correctly, as they can learn the training data using different combinations of conjunctions. Notably, kernel models trained with gradient descent or ridge regression generally learn the readout weights with minimal $\ell_{2}$-norm [[, Section 2;]]soudry_implicit_2018,gunasekar_characterizing_2018,ji_gradient_2020. This inductive bias gives rise to weights that are distributed across different predictive features. For in-distribution generalization, this is often useful, as it allows us to integrate different kinds of evidence. However, for compositional generalization, we often need to rely on a small group of conjunctions and the tendency towards distributed weights causes generalization failures.
Here we highlight two particular failure modes arising from this. First, the full conjunctions always serve as predictive features (“memorization leak”). Second, the model can often use statistical shortcuts to predict parts of the training set, using the full conjunction to learn the rest (“shortcut bias”). Below we explain how these failure modes impact symbolic addition and context dependence.
Symbolic addition suffers from a memorization leak. The functional form $f(z)=f_{1}(z_{1})+f_{2}(z_{2})+f_{12}(z)$ implies that using the full conjunction $f_{12}(z)$ to learn the training set will necessarily distort the inferred values $f_{1}(z_{1}),f_{2}(z_{2})$. This, in turn, negatively impacts generalization on the test set. To analyze the impact of such a memorization leak in detail, we characterize model behavior on symbolic addition analytically for training sets that are balanced around zero:
Proposition 4.4.
Consider input elements in $\{[v]\}_{v\in\mathcal{V}},\mathcal{V}\subset\mathbb{R}$ with associated values $v$. We assume that the training set contains all pairs such that at least one component is $z_{c}\in\{[w]\}_{w\in\mathcal{W}},\mathcal{W}\subset\mathbb{R}$ and that the average value in both $\mathcal{V}$ and $\mathcal{W}$ is zero. Then, model behavior on the test set is given by
$f([i],[j])=m(i+j),\quad m:=\tfrac{p\cdot\text{Sal}_{2}(1)}{1+(p-2)\text{Sal}_{2}(1)},\quad p:=|\mathcal{W}|$ (7)
We prove the proposition in Section D.1.1. It implies that for any representation that encodes the conjunction (i.e. has $\text{Sal}_{2}(2)>0$ and therefore $\text{Sal}_{2}(1)<\tfrac{1}{2}$), the model underestimates the inferred values by a constant factor $m<1$. Further, $m$ decreases for smaller $\text{Sal}_{2}(1)$ and increases for larger training set size (Fig. 3a,b). Perhaps surprisingly, these are the only two factors influencing $m$. In particular, while we often assume that interpolation is easier than extrapolation, these changes in training set do not impact model behavior here. We found qualitatively similar behavior for randomly generated, dispersed training sets (Section D.1.2, Fig. 7).
Figure 3: Kernel models’ behavior on (a,b) symbolic addition and (c,d) context dependence. a, Factor $m$ distorting the inferred values as a function of $\text{Sal}_{2}(1)$ and the training set size $p=|\mathcal{W}|$. b, Inferred values on extrapolation task. c, Variants of context dependence for which different representational saliences yield successful generalization. Trajectories of networks with different nonlinearities are highlighted (color scale see Fig. 2). d, Inferred coefficients of the different conjunction types for two example representations: a network with three layers and a rectified linear or quadratic nonlinearity.
Notably, because the full conjunction can always be used to learn the training set, the memorization leak impacts a broad range of compositional tasks, including partial exposure (Section D.4).
Context dependence suffers from a shortcut bias. Next, we empirically analyzed model generalization on context dependence across different representational geometries and training sets. We found that on a given task, each model either generalizes with 100% or 0% accuracy. For CD-3, the model only generalizes when $\text{Sal}_{3}(2)$ is high relative to $\text{Sal}_{3}(1)$ (Fig. 3c). As a result, whether the network generalizes is highly sensitive to the nonlinearity and depth of the network. In contrast, for CD-2 and CD-1, a much wider range of representational geometries generalizes successfully.
To better understand this, we analyzed the total magnitude of model weights associated with the different conjunctions. Notably, there are many possible conjunctions the model could rely on (see Section D.2.4). In particular, context $1$ is more likely to have the target $y=1$ and context $2$ is more likely to have target $y=-1$. The model may thus rely on this statistical shortcut, using the full conjunction to memorize the remaining training data. Indeed, we found that representations that generalized unsuccessfully on CD-3 consistently had a high weight associated with the context component and the full conjunction (Fig. 3d). Notably, for CD-3, the context shortcut yields an accuracy of $\tfrac{2}{3}$. This explains why when $\text{Sal}_{3}(1)$ and $\text{Sal}_{3}(3)$ are high relative to $\text{Sal}_{3}(2)$, the model uses the context-driven shortcut together with the full conjunction, resulting in a failure to generalize. For CD-2, the context-driven shortcut only results in an accuracy of $\tfrac{9}{16}$, making it much less useful. This explains why only very low $\text{Sal}_{3}(2)$ yields failure to generalize on CD-2 or CD-1.
Our analysis illustrates how conjunction-wise additivity lets us identify component conjunctions that yield a shortcut. The representational salience of these conjunctions and the extent to which they can explain the training set then helps us understand when this shortcut impacts generalization.
4.4 Conjunction-wise additivity can describe generalization behavior in deep neural networks
So far, we have analyzed kernel models, for their analytical tractability. We conjecture that large-scale neural network architectures with more complex, interrelated inputs might also be prone to implementing a conjunction-wise additive computation. In that case, our analyis of the kernel models would shed light on when the statistics of the training data enable or prevent compositional generalization in deep neural networks as well. To test our conjecture, we trained convolutional networks (ConvNets, [lecun_backpropagation_1989]) on a version of the tasks with concatenated MNIST digits [lecun_gradient-based_1998] and residual networks (ResNets, [he_deep_2016]) and Vision Transformers (ViTs, [dosovitskiy_image_2020]) on a version of the tasks with concatenated CIFAR-10 images [krizhevsky_learning_2009]. We considered the same tasks except that each component was now given by images of a particular category, rather than being a single instance. Notably, different image categories are not necessarily equally correlated with each other. To control for this, we randomly permuted the assignment of categories to components for each ($n=10$) experiment. To compare the network behaviors to our theoretical predictions, we fit a conjunction-wise additive function to the network output using linear regression (“additivity analysis”; see Section C.3).
Symbolic addition. We trained the ConvNets on a total of 20,000 randomly generated MNIST samples for 100 epochs and trained the ResNets and ViTs on a total of 40,000 CIFAR-10 samples for 100 and 200 epochs, respectively. Remarkably, we found that the kernel theory matches network behavior almost perfectly: the networks’ average predictions (across the different possible permutations of the digit categories) were well predicted by an additive structure and their inferred values generally underestimated the ground truth (Fig. 4a, Figs. 9 and 10). This suggests that the kernel theory can capture the behavior of relevant neural network architectures trained on natural image inputs.
Next, to investigate the effect of increased conjunctivity, we varied the distance between the two MNIST digits. We hypothesized that the ConvNets’ local weight structure should produce a more conjunctive representation for digits that are closer together. To test this, we determined $\text{Sal}_{2}(1)$ in an intermediate layer of the network, averaging over different instances of all digits. We found that $\text{Sal}_{2}(1)$ was indeed smaller for lower distances (Fig. 4b). Further, the inferred values were more compressed for smaller distances, confirming that more conjunctive inputs exacerbate the memorization leak in ConvNets as well (Fig. 4c and Fig. 9).
Figure 4: Behavior of networks trained on MNIST and CIFAR versions of compositional tasks. Shaded regions indicate mean $\pm$ two standard errors. a, Inferred values on symbolic addition (extrapolation, distance of zero between digits). b, $\text{Sal}_{2}(1)$ in an intermediate layer of the ConvNet, for different distances between digits. c, Inferred values for the extrapolation task and different distances between digits. d, For the networks trained on variants of context dependence, the inferred coefficients for different conjunction types (distance of 20px).
Context dependence. We then trained the ConvNets on an MNIST version of context dependence using 30,000 training samples and the ResNets and ViTs on a CIFAR-10 version of the task using 40,000 training samples. Again, their behavior was aligned with the kernel theory’s predictions, having better-than-chance accuracy on CD-1 and CD-2, but failing to generalize on CD-3. The additivity analysis revealed that this was due to a context-driven shortcut (Fig. 4d).
4.5 Feature learning can overcome the limitations of kernel models
We proved above that compositionally structured kernel models do not generalize on non-additive tasks, including transitive equivalence. To see if feature learning can overcome this limitation, we trained ReLU networks through backpropagation on transitive equivalence, using a disentangled and uncorrelated input. By varying the initial weight magnitude, we either trained these networks in the kernel/lazy regime or the feature-learning/rich regime. Notably, when trained on symbolic addition and context dependence, the rich networks were well-described by a conjunction-wise additive model (Figs. 8 and 11). On transitive equivalence, however, while the kernel-regime models failed to generalize (as predicted by our theory), the rich networks generalized correctly (Fig. 5a).
To explain why this is the case, we leveraged the insight that rich neural networks are biased to learning weights with a low overall $\ell_{2}$-norm ([lyu_gradient_2020, chizat_implicit_2020]; cf. [vardi_implicit_2021]). In particular, a one-hidden-layer ReLU network tends to learn a sparse set of features [savarese_how_2019, chizat_implicit_2020]. Transitive equivalence consists of multiple overlapping equality relations (e.g. $A=B\neq D=E$). Notably, ReLU networks such an XOR-type problem by specializing one unit to each conjunction ([brutzkus_why_2019, saxe_neural_2022, xu_benign_2023]; Fig. 5b). Further, their sparse inductive bias incentivizes ReLU networks to use identical units for overlapping conjunctions (e.g. $(A,B),(A,C)$, and $(B,C)$). This causes the unit to generalize to unseen item combinations (e.g. $(A,C)$), enabling the network to generalize. Importantly, our theoretical argument is corroborated by empirical simulations: each network unit has identical weights for equivalent items (Fig. 5c).
To see whether large-scale neural networks can also benefit from this feature-learning mechanism, we trained ConvNets on an MNIST version of transitive equivalence. The networks were trained for 150 epochs on 20,000 samples. We found that if the digits were presented with a distance of zero, the network did not generalize compositionally at all. However, with increasing distance, the network started to improve its compositional generalization (Fig. 5e), demonstrating that a convolutional network can benefit from this rich compositional motif.
5 Discussion
Humans often generalize to new situations by stitching together concepts and knowledge from prior experience. Despite the broad importance of this ability (both for cognitive science and machine learning), it has remained unclear under what conditions compositionally structured representations give rise to compositional generalization. Here we have taken a step toward formalizing this relationship by clarifying the full range of compositional motifs implemented by kernel models with categorical compositionally structured representations (“conjunction-wise additivity”). For conjunction-wise additive tasks, we analyzed how representational geometry and training data impact successful generalization on important compositional building blocks, validating our analysis for deep neural networks trained on natural image data. For non-additive tasks, our results immediately imply that kernel models are unable to generalize. We then analyzed how feature-learning mechanisms can overcome these limitations and demonstrated that deep networks benefit from the same mechanisms. Taken together, this suggests that conjunction-wise additivity provides a useful generalization class for understanding compositional generalization in neural networks and perhaps even humans.
Limitations. We considered tasks with a fixed number of input components and fixed output dimensionality and assumed that all input components are represented orthogonally, excluding continuously varying components such as position. Further, our theoretical analysis focused on kernel models with fixed representations (though we considered deep neural networks through empirical simulations). Future work could extend the presented theory beyond these assumptions, and could further explore the range of compositional motifs in other learning models (e.g. rich neural networks and modular networks) more systematically. Nevertheless, we believe that this work is an important step towards a comprehensive theory of compositional generalization, covering a broad range of tasks and highlighting several phenomena that might help us understand more complex models as well.
Acknowledgments
[…]
Appendix A Proof of Theorem 4.1
@@ Acknowledgments
Because $K$ is compositionally structured, its similarity $K(I,I^{\prime})$ only depends on the overlap $O(I,I^{\prime})\subseteq\{1,\dotsc,C\}$. We denote the similarity for inputs overlapping in $S\subseteq\{1,\dotsc,C\}$ by $\kappa_{S}$ and define the overlap in the training dataset as
$\mathcal{D}_{S}(I):=\left\{I^{\prime}\in\mathcal{D}^{(\text{train})}|\forall_{c\in S}I_{c}=I^{\prime}_{c}\right\}.$ (8)
The key idea is to decompose $\mathcal{D}^{(\text{train})}$ into these different overlaps in order to separate the sum into its components. However, by our definition, the datasets $\mathcal{D}_{S}(I)$ are not disjoint. Indeed, $S\subseteq S^{\prime}$ implies $\mathcal{D}_{S^{\prime}}(I)\subseteq\mathcal{D}_{S}(I)$ and in particular $\mathcal{D}_{\emptyset}(I)=\mathcal{D}^{(\text{train})}$. To adjust for this, we define $\delta_{S}$ as the similarity added by $\kappa_{S}$ to the similarity between conjunctions with one component fewer, recursively defining
$\delta_{\emptyset}=\kappa_{\emptyset},\quad\delta_{S}=\kappa_{S}-\sum_{S^{\prime}\subsetneq S}\delta_{S^{\prime}}.$ (9)
@@
$f(I)=\sum_{I^{\prime}\in\mathcal{D}^{(\text{train})}}a_{I^{\prime}}K(I,I^{\prime})=\sum_{S\subseteq\{1,\dotsc,C\}}\delta_{S}\sum_{I^{\prime}\in\mathcal{D}_{S}(I)}a_{I^{\prime}}.$ (10)
This equality obtains because for each $I^{\prime}\in\mathcal{D}^{(\text{train})}$,
$\sum_{S:I^{\prime}\in\mathcal{D}_{S}(I)}\delta_{S}=\delta_{O(I,I^{\prime})}+\sum_{S\subsetneq O(I,I^{\prime})}\delta_{S^{\prime}}=\kappa_{O(I,I^{\prime})}=K(I,I^{\prime}),$ (11)
which is true by definition. We note that for $S\notin\mathcal{O}(I)$, $\mathcal{D}_{S}(I)=\emptyset$. Defining
$f_{S}(I):=\delta_{S}\sum_{I^{\prime}\in\mathcal{D}_{S}(I)}a_{I^{\prime}},$ (12)
@@
Appendix B Representational geometries across different deep neural network architectures
We computed the saliences by iteratively computing the representational similarities using the kernels derived in prior work [cho_kernel_2009, tsuchida_invariance_2018, tsuchida_richer_2019, han_fast_2022].
B.1 Proof of Proposition 4.3
See 4.3

### v2

@@ A.3 Nonlinear transformations and compositional structure
The infinite-width limit is characterized by $H_{1},\dotsc,H_{L-1}\to\infty$ (the order of these limits does not matter). This setup includes, in particular, the random feature model investigated by Abbe et al. (2023).
Further, the condition is met by most commonly used kernel functions including any radial basis function that depends on the Euclidean $\ell_{2}$-norm (e.g. the Gaussian kernel). Any nonlinear transform that meets this condition also conserves compositional structure, i.e. if $z$ is compositionally structured then so is $\phi(z)$. This allows us to derive computational restrictions on this broad range of models.
A.4 Proof of Theorem 4.1
@@
Because the kernel $K$ is compositionally structured, its similarity $K(z,z^{\prime})$ only depends on the overlap $O(z,z^{\prime})\subseteq\{1,\dotsc,C\}$ (see definition in Eq. 6). We denote the similarity for inputs overlapping in $S\subseteq\{1,\dotsc,C\}$ by $\kappa_{S}$ and define set of training items overlapping with $z\in Z^{\text{test}}$ in $S$ as
$Z^{\text{train}}(z,J):=\left\{z^{\text{tr}}\in Z^{\text{train}}|\forall_{c\in J}z_{c}=z^{\text{tr}}_{c}\right\}.$ (8)
The key idea is to decompose $Z^{\text{train}}$ into these different overlaps in order to separate the sum into its components. However, by our definition, the datasets $Z^{\text{train}}(z,J)$ are not disjoint. Indeed, $S\subseteq S^{\prime}$ implies $Z^{\text{train}}(z,J)\subseteq Z^{\text{train}}(z,J^{\prime})$ and in particular $Z^{\text{train}}(z,\emptyset)=Z^{\text{train}}$. To adjust for this, we define $\delta_{J}$ as the similarity added by $\kappa_{S}$ to the similarity between conjunctions with one component fewer, recursively defining
$\delta_{\emptyset}=\kappa_{\emptyset},\quad\delta_{J}=\kappa_{J}-\sum_{J^{\prime}\subsetneq J}\delta_{J^{\prime}}.$ (9)
@@
$f(z)=\sum_{z^{\text{tr}}\in Z^{\text{train}}}a_{z^{\text{tr}}}K(z,z^{\text{tr}})=\sum_{J\subseteq\{1,\dotsc,C\}}\delta_{S}\sum_{z^{\text{tr}}\in Z^{\text{train}}(z,J)}a_{z^{\text{tr}}}.$ (10)
This equality obtains because for each $z\in Z,z^{\text{tr}}\in Z^{\text{train}}$,
$\sum_{J:z^{\text{tr}}\in Z^{\text{train}}(z,J)}\delta_{J}=\delta_{O(z,z^{\text{tr}})}+\sum_{J^{\prime}\subsetneq O(z,z^{\text{tr}})}\delta_{J^{\prime}}=\kappa_{O(z,z^{\text{tr}})}=K(z,z^{\text{tr}}),$ (11)
which is true by definition. We note that for $J\notin\text{Conj}(z|Z^{\text{train}})$, $Z^{\text{train}}(z,J)=\emptyset$. Defining
$f_{J}(z):=\delta_{J}\sum_{z^{\text{tr}}\in Z^{\text{train}}(z,J)}a_{z^{\text{tr}}},$ (12)
@@
Appendix B Representation salience: a metric for compositionally structured representations
B.1 Definition
Below we formally define representational salience. To do so, we denote the similarity between two trials $z,z^{\prime}$ by $\text{Sim}(O(z,z^{\prime})):=K(z,z^{\prime})$. Note that this is well-defined for compositionally structured representations, as $K(z,z^{\prime})$ is identical for all pairs of trials with the same $O(z,z^{\prime})$. We then define:
Definition B.1.
For a generalized compositionally structured representation with kernel $K$ and a conjunction $J\subseteq\{1,\dotsc,C\}$, we recursively define
$\overline{S}(\emptyset):=K(\emptyset),\quad\overline{S}(J):=\text{Sim}(J)-\sum_{J^{\prime}\subsetneq J}\overline{S}(J^{\prime}),\quad S(J)=\tfrac{\overline{S}(J)}{\sum_{\emptyset\neq J\subseteq\{1,\dotsc,C\}}\overline{S}(J)}.$ (13)
When emphasizing the total number of components, we denote $S(J;C)=S(J)$. Further, for compositionally structured representation, the similarity only depends on the total number of overlaps $|J|$. We therefore write $S(c;C)=S(J)$, where $c:=|J|$.
B.2 Why is representational salience a useful metric?
Intuitively, the representational salience captures the unique contribution of a population representing a particular conjunction $J$ (as in our informal definition in the main text). As we noted in the main text, distinguishing these unique contributions from the similarities directly would be more difficult. For example, changes in the similarity between inputs having two components in common could arise from changes in how saliently single components or conjunctions of two components are represented. To distinguish between these cases, we’d have to additionally look at the similarity between inputs having a single component in common. Our definition of representational salience avoids this issue; we therefore believe that this perspective could more broadly be useful for analyzing representations of compositional data.
B.3 Why does representational salience leave out magnitude and baseline activity?
Notably, many learning models are largely invariant to a constant rescaling of their representation. For limit behavior (e.g. in the limit of infinite training in gradient descent), this scale has no impact. We see this, for instance, in Proposition 5.2, where the scaling turns out to be entirely irrelevant. For regularized models, on the other hand, the scale of the representation is confounded with the strength of the regularization and so we suggest that it is again best seen as separate from the representational geometry.
The baseline activity, on the other hand, may determine how easily an intercept is learned. However, many linear readout models (e.g. support vector machines) do not regularize their intercept, again rendering this parameter entirely irrelevant. Notably, in the case of gradient descent, the magnitude of the baseline activity does influence how easily an intercept is learned. However, the impact of this is often negligible; for example, in Proposition 5.2, we again find that the baseline activity is entirely irrelevant.
B.4 Extended representational analysis
We computed the saliences by iteratively computing the representational similarities using the kernels derived in prior work (Cho & Saul, 2009; Tsuchida et al., 2018, 2019; Han et al., 2022). In Fig. 5b, we plot the different saliences for an input with four components. Now the salience of overlap 2 and 3 both first increase and then decrease and, just like for inputs with three components (Fig. 2), a Gaussian and rectified quadratic nonlinearity yields a particularly high salience for these intermediate conjunctions. Notably, they appear to be trading off the salience of these conjunctions differently: the rectified quadratic nonlinearity more strongly emphasizes overlaps of three whereas the Gaussian nonlinearity more strongly emphasizes overlaps of two. This highlights that for larger numbers of components, the dependence of generalization behavior on specific architectural choices will likely be even stronger.
Finally, we consider an example of a generalized compositionally structured representation. Here we assume that the different input components have different magnitudes. In particular, we consider a disentangled representation whose first component has a salience between $s\in[0.1,0.5]$ and whose second component accordingly has a salience $1-s$ (Fig. 5a). As the random weight neural network becomes deeper, the more salient component (in this case component 2) increasingly dominates the representation. While the salience of the full conjunction still eventually converges to one for most nonlinearities, it takes longer to do so for a less balanced input representation. Further, for a Gaussian nonlinearity, an imbalanced representation actually decreases the limit salience the representation appears to be converging to for the full conjunction. This highlights that for disentangled representations that do not represent all components with equal magnitude, compositional generalization behavior may vary strongly with different neural network architectures.
B.5 Proof of Proposition 5.1
See 5.1